OpenStudy (anonymous):

Given that P(A) = 0.3, P(B) = 0.6, and A and B are independent events, compute each probability. (b) P(A or B)

OpenStudy (anonymous):

P(A)+P(B)-P(A)x P(B) = 0.3+0.6-0.3 x 0.6 = 0.72